OpenStudy (jessiegonzales):

MEDAL!!!! A spherical balloon has a circumference of 19cm. What is the approximate surface area of the balloon to the nearest square centimeter? A. 115cm^2 B. 460cm^2 C.361cm^2 D.1,134cm^2

OpenStudy (anonymous):

find radius frm circumference

OpenStudy (fibonaccichick666):

so first, what is the formula for surface area?

OpenStudy (jessiegonzales):

what is the formula again?

OpenStudy (fibonaccichick666):

well I am asking you, jessie

OpenStudy (anonymous):

\[4*\pi *r*r=surface area of sphere \]

OpenStudy (fibonaccichick666):

ok, so jessie, now what is the formula for circumference

OpenStudy (jessiegonzales):

4*pi*r*r

OpenStudy (fibonaccichick666):

that is not circumference 4pi r squared is the formula for surface area

OpenStudy (jessiegonzales):

so whats the formula i forgot

OpenStudy (anonymous):

circumference = 2*pi*R

OpenStudy (jessiegonzales):

2*pi*9.5 right

OpenStudy (jessiegonzales):

59.66

OpenStudy (fibonaccichick666):

you don't know r yet, so you cannot put it in. You need to solve \[C=2\pi r\] for r. Please do this

OpenStudy (jessiegonzales):

ok.. 19=2pi r

OpenStudy (anonymous):

A)

OpenStudy (jessiegonzales):

how did u get A

OpenStudy (fibonaccichick666):

so solve that for r now please